# Postmerger: a new and dominant contribution to the gravitational-wave background from binary neutron stars

### v1

@@ Abstract
The stochastic gravitational-wave background (SGWB) generated by the inspiral and merger of binary neutron stars is traditionally modelled assuming that the inspiral is promptly followed by the collapse of the merger remnant to a rotating black hole. While this is reasonable for the most massive binaries, it is not what is expected in general, where a remnant is produced and may survive for up to hundreds of milliseconds and radiate an amount of energy that is significantly larger than that lost during the whole inspiral. To account for this additional contribution to the SGWB, we consider a waveform model that includes both the inspiral and the postmerger emission. We show for the first time that for a large set of parameterized equations of state compatible with observational constraints, there is typically five times more power spectral density in the SGWB from the postmerger emission than from the inspiral one, leading to a normalized GW energy density $\Omega_{\rm GW}\simeq 10^{-10}-10^{-9}$. This power is predominantly located in the $1-2\,{\rm kHz}$ frequency range, hence distinct from that associated with the inspiral only. We discuss the significantly enhanced detectability of the SGWB with special attention to third-generation detectors, such as the Einstein Telescope and Cosmic Explorer, and show how it depends on the signal-to-noise ratio of foreground binaries and on the metastable remnant survival time. Interestingly, even a non-detection of the high-frequency part of the SGWB could provide valuable constraints on the remnant lifetime, offering novel insights into the postmerger dynamics and into the equation of state of nuclear matter.
I Introduction
@@
The discovery of the binary neutron star (BNS) merger GW170817 [1] marked a major milestone in multi-messenger astrophysics, as the observations of gravitational waves (GW) by LIGO-Virgo [2, 3] was followed by the detection of electromagnetic counterparts across multiple wavelengths [4, 5, 6, 7, 8]. These observations confirmed the connection between short gamma-ray bursts, kilonovae and BNS mergers and provided important information on the geometry of the jet launched after the merger as well as on the ejected neutron-rich material. Joint analysis of electromagnetic and GW data led to tight constraints of General Relativity [9], whereas the identification of the host galaxy allowed an independent measurement of the Hubble parameter (see, e.g., [10]). Combined with the second, GW-only, BNS detection [11], the local merger rate is constrained to $10-1700\,{\rm Gpc}^{-3}\,{\rm yr}^{-2}$ [12]. This information is crucial for building population models and understanding the formation of binary compact objects.
@@
Observations of BNS mergers also provide a unique probe of the physics of extremely dense matter and allow to study the equations of state (EOS) of neutron star interiors. Important constraints on the EOSs can come from electromagnetic observations of isolated neutron stars or via GW observations of BNS mergers, as demonstrated with the detection GW170817. Since the EOS strongly affects the gravitational waveform, it leaves an imprint in the tidal deformation takes place in the late inspiral stages and allows measuring the tidal deformability [14, 15, 16, 17, 18]. In the case of GW170817, GW observations allowed to constrain the tidal deformability parameter to $\tilde{\Lambda}\lesssim 700$, where $\tilde{\Lambda}$ is the mass-weighted combination of the dimensionless deformability parameters [19, 20].
@@
Additional and complementary information on the EOS can be extracted from the postmerger signal emitted by the hypermassive neutron star (HMNS) formed after the neutron stars have coalesced [21, 22, 23]. In particular, a number of recent studies have shown that the GW emission from the HMNS is concentrated at specific frequencies, resulting in relatively narrow spectral features [24, 25, 26, 27, 28, 29, 30]. The postmerger signal has not been detected so far, as the relevant frequencies are above those observable with current ground-based detectors. However, future third-generation (3G) detectors, such as the Einstein Telescope (ET) [31] and the Cosmic Explorer (CE) [CE: 32] will be able to detect and characterize the postmerger signal for sufficiently nearby systems.
A major limitation of GW observations of BNS mergers is that they are difficult to observe at large distances. Given typical BNS masses, current detectors are limited to only $\sim 500\,{\rm Mpc}$ [33]. While ET and CE will be able to reach $z\sim 2-3$ [34], in order to accurately measure the postmerger signal a high signal-to-noise ratio (SNR) is required, which is only achievable for relatively nearby systems (up to a few tens of Mpc) [35].
@@
An alternative method to study populations of compact binaries is by observing the stochastic GW background (SGWB) created by the incoherent superposition of unresolved sources [36, 37]. Different types of sources can contribute to the SGWB, including compact binaries, core-collapse supernovae and isolated spinning neutron stars [e.g. 38, 39, 40]. Observing the SGWB will allow probing high-redshift populations, which are otherwise not accessible to ground-based detectors, although various methods are being developed to extract the SGWB from the detector noise in data from Advanced LIGO, Advanced Virgo and Kagra (LVK) are ongoing [41, 42]. Current upper limits for ground-based detectors are $\Omega_{\rm{GW}}(f)\leq 3.4\times 10^{-9}$ at $f=25$ Hz, where $\Omega_{\rm{GW}}$ is the dimensionless energy density of the SGWB [43, 44, 45, 46, 47]. These upper limits contain valuable information on high-redshift populations of compact binaries, complementary to that of individual detections [e.g. 48, 49, 50, 51, 52, 53, 54, 55]. For the LVK detectors, the dominant contribution to the SGWB is expected to be that of binary black holes. However, this will not be the case for ET and CE, since almost all the binary black hole mergers will be detected individually and subtracted from the data. The remaining SGWB will thus contain information on the unresolved BNS mergers.
@@
In the case of the postmerger emission from a population of BNS mergers, we expect that if HMNS are formed, the characteristic GW frequencies will be enhanced in the SGWB due to the large number of unresolved sources contributing to the signal. This opens up the possibility to detect this EOS-dependent feature and to obtain complementary information on the postmerger stage. The goal of this study is to model the SGWB from a population of BNS, including the postmerger stage, and to evaluate its detectability with ET and CE. This paper is structured as follows. In Section II we present the BNS waveform used in this study, our method to construct a distribution of EOSs and discuss the calculation of the BNS background. We present our results in Section III and discuss future prospects in Section IV.
Figure 1: Evolution of the radiated GW energy when normalized to the at the merger. Shown with blue- and red-shaded areas are the corresponding values during the inspiral (blue) and the postmerger. The variance in the shaded regions marks the $99\%$ percentile and reflects the range of losses associated with the population of BNS considered here. Note that the postmerger energy can easily be $3-4$ larger than the inspiral one over the first $\sim 20\,{\rm ms}$.
II Methodology
II.1 BNS inspiral and postmerger waveform
A major focus of our work is to extend the contribution of the BNS merger signal to the SGWB by considering not only the inspiral part but also the postmerger emission. To appreciate simply why this is important and hence the relevance of our findings, we show in Fig. 1 the total GW energy radiated by a BNS when integrated over the entire frequency range ($E_{\mathrm{GW}}$) and normalized by the total energy considering only the inspiral phase ($E_{\mathrm{GW},0}$), as a function of the time before and after the merger $t-t_{\rm mer}$. Near the merger ($t-t_{\rm mer}\lesssim 0$) the energy radiated increases very rapidly, gaining an order of magnitude in the last $10\,{\rm ms}$ before the merger. Following the merger ($t-t_{\rm mer}>0$), the energy radiated continues to grow at a slower rate but more intensely, so that it can reach values much larger than those lost during the whole inspiral. Stated differently, BNS mergers are copious emitters of GWs during the inspiral, but it is the postmerger that provides the largest contribution to the radiated energy.
Of course, the exact amount of energy radiated will depend sensitively on the EOS describing the matter of which the BNS is composed of and which is essentially unknown. Instead of concentrating on a few specific EOSs, it is more convenient to consider a large ensemble of agnostic EOSs satisfy astrophysical, GW, or elementary-particle constraints [56, 57, 58, 59]. This is what is represented with the blue-shaded band in Fig. 1, which reports the 99th percentile contour of an ensemble of $10^{3}$ EOSs generated using our EOS-agnostic approach (see Sec. II.2). Similarly, the solid blue line reports the median of the distribution at each time in the evolution. In this way, it is straightforward to appreciate that since most of the energy is lost after the merger, it is important to include also the postmerger emission when calculating the BNS GW background (similar considerations apply when considering the radiated angular momentum [60, 61]).
@@ II.1 BNS inspiral and postmerger waveform
To construct a hybrid waveform, we combine in the frequency domain the analytic prescriptions referring to the inspiral and the postmerger. As presented in detail in Appendix A, this combination takes place at a frequency $f_{\rm tr}$, where we move from the inspiral (see Sec. A.1) to the postmerger (see Sec. A.2) description. This frequency must be significantly smaller than the main frequencies appearing in the power spectral density (PSD) of the postmerger signal, i.e., $f_{1}$ and $f_{2}$, where $f_{1}<f_{2}$ (see Sec. A.2) so that no power from the postmerger phase is missed. Hereafter, for simplicity we set $f_{\rm tr}=f_{1}/2$, but the overall results depend only weakly on this assumption.
Figure 2 illustrates the construction of the hybrid waveform for a BNS system with masses $M_{1}=M_{2}=1.25\,M_{\odot}$, at a distance from observer $D=50\,{\rm Mpc}$ when employing GNH3 EOS discussed in Ref. [62]. Shown with a blue dashed line is the inspiral-and-merger signal followed by a prompt collapse and described by the IMRPhenomD_NRTidalv2 model, the postmerger signal with a black dashed line, and the combined (hybrid) signal with a red solid line. We also show with a solid green line the ET strain sensitivity. The transition frequency $f_{\rm tr}$ is meant to mark therefore the boundary between the grey and white regions. In the grey region, the hybrid waveform is simply given by the IMRPhenomD_NRTidalv2 waveform. Note that in the white region, the hybrid waveform is given by the sum of the inspiral and postmerger strains and that the latter is negligible below $f_{\rm tr}$, thus ensuring no significant power is lost or added. To properly combine the two waveforms above $f_{\rm tr}$, we must ensure phase coherence, which is not guaranteed a priori, as we have not explicitly computed the temporal phase $\beta_{2}$ in Eq. (8). Therefore, at $f_{\rm tr}$, we adjust the phase of the IMRPhenomD_NRTidalv2 waveform to match the postmerger one, thus ensuring phase continuity.
Figure 2: Evolution in the frequency-domain of the GW strain for a BNS with masses $m_{1}=m_{2}=1.25\,M_{\odot}$ at a distance of observer $D=50\,{\rm Mpc}$ and relative to the GNH3 EOS [62]. The IMRPhenomD_NRTidalv2 waveform is shown with a black dashed line, the HMNS postmerger with a blue dashed line, and the combined waveform with a red solid line. The frequency $f_{\rm tr}$ marks the transition between the inspiral (grey-shaded are) and the postmerger, while the green solid curve reports the sensitivity of ET
II.2 Modelling the equations of state
In view of the uncertainties in the EOS and their effect on the waveform, we adopt an agnostic approach in this work. We use the findings of Refs. [59, 63, 64, 65], who constructed more than $10^{7}$ EOSs consistent with constraints from fundamental physics and astronomical observations of neutron stars. Leveraging on this extensive dataset, it was possible to derive informative constraints on the relationship between the chirp mass $\mathcal{M}_{\mathrm{chirp}}$ and tidal deformability $\Lambda$ in terms of a probability density function (PDF) for these two quantities. We have exploited this correlation to create a consistent and agnostic description, assigning a range of allowed values of the tidal deformability for any choice of the chirp mass.
@@ II.2 Modelling the equations of state
A-priori, a single EOS does not correspond to a unique line in the $\mathcal{M}_{\mathrm{chirp}}$-$\Lambda$ plane because the same chirp mass may result from many combinations of BNS masses. However, when restricting the analysis to symmetric binaries, as we do throughout the rest of this work, this ambiguity disappears. In this case, for a given EOS, each chirp mass is uniquely associated to a single value of $\Lambda$. Therefore, extracting an EOS from a given distribution is equivalent to drawing a single value of $\Lambda$ from the PDF at each chirp mass. Similarly, extracting a certain value of $\Lambda$ from the allowed range is equivalent to singling-out a specific EOS.
Although the PDF found in Ref. [59] for a given chirp mass is not exactly a Gaussian, it is so to a first approximation [see Fig. (4) in [59]]. Hence, when assuming an underlining Gaussian distribution, we can relate the tidal deformability and the chirp mass (in solar masses) via a simple power-law of the type
$\Lambda_{p}=\alpha+\beta\mathcal{M}^{\gamma}_{\mathrm{chirp}}\,,$ (1)
where $\Lambda_{p}$ is normally distributed between the minimum and maximum values given by [59]
$\Lambda_{\rm min\,(max)}=a+b\,\mathcal{M}_{\rm chirp}^{c}\,,$ (2)
with $a=-50\,(-20)$, $b=500\,(1800)$, and $c=-4.5\,(-5.0)$. In this way, we can build a set of fictitious $10^{3}$ EOSs such that the corresponding tidal deformability follows expression (1) and is normally distributed.
@@
Once an EOS is set, we can compute the inspiral waveform as discussed in Sec. A.1. For the postmerger waveform, and as discussed in Sec. A.2, we need to know the characteristic frequencies $f_{1}$ and $f_{2}$ given the selected EOS and hence tidal deformability. Luckily, this is rather straightforward to do since universal relations have been found that provide analytic fits (see, e.g., [25, 26, 27, 28]). Here, we employ the expressions proposed in Refs. [28, 30] (and also employed in [62]) and express the postmerger frequencies $f_{1}$ and $f_{2}$ as
$\frac{f_{1}}{\mathrm{kHz}}$ $=$ $c_{0}+c_{1}x+c_{2}x^{2}+c_{3}x^{3}\,,$ (3)
$\frac{f_{2}}{\mathrm{kHz}}$ $=$ $10^{\,a_{0}+b_{0}\,x}\,\left(\frac{M_{\odot}}{M}\right)\,,$ (4)
where $x:=(3\Lambda/16)^{1/5}$, $M:=m_{1}+m_{2}$ is the total mass of the system, and the constants are $c_{0}=45.195$, $c_{1}=-43.484$, $c_{2}=14.653$, $c_{3}=-1.662$, $a_{0}=1.617$, and $b_{0}=-0.274$.
We conclude this section by remarking that expressions (4), being only “quasi”-universal relations, obviously come with some intrinsic uncertainties as estimated from the numerical simulations and the corresponding fits (see discussion in [30]). These uncertainties, however, do not play an important role here as we are interested in a SGWB that is naturally broad given the much larger uncertainties in the EOS.
II.3 Stochastic gravitational-wave background from BNS mergers
The energy density in GWs per logarithmic unit of frequency and in units of the critical density of the Universe, $\rho_{\mathrm{c}}=3H_{0}/8\pi G$, is given by
$\Omega_{\mathrm{GW}}:=\frac{1}{\rho_{c}}\frac{\mathrm{d}\rho_{\mathrm{GW}}}{\mathrm{d}\ln{f}}=\frac{f}{c\,\rho_{\mathrm{c}}}F(f)\,,$ (5)
where $\rho_{\mathrm{GW}}$ is the GW energy density and $F(f)$ is the energy flux. When considering a population of $N$ sources, the energy flux per unit frequency is given by [66]:
$F(f):=\frac{\pi c^{3}}{2G}\frac{f^{2}}{T}\sum_{i=1}^{N}\left[|\tilde{h}_{i}^{+}(f)|^{2}+|\tilde{h}_{i}^{\times}(f)|^{2}\right]\,,$ (6)
@@ II.3 Stochastic gravitational-wave background from BNS mergers
where the index $i$ sums over the $N$ sources and $\tilde{h}_{i}^{+}(f)$ and $\tilde{h}_{i}^{\times}(f)$ are the Fourier transforms of the two GW polarization modes. The number of sources is taken to be proportional to the total observation time $T$, which in this study is set to one year. We have checked that our calculation converges at $T=1$ year and will use this value in the following (note that the background is stationary).
In order to construct our BNS catalogue, we follow the baseline_delays model from Ref. [55] and, more specifically, we assume all binaries are symmetric in mass and follow a Gaussian mass distribution with a mean of $1.33\,M_{\odot}$ and a standard deviation of $0.09\,M_{\odot}$. This distribution is consistent with the theoretical upper limits on the maximum mass of a nonrotating NS of $\sim 2.16-2.3M_{\odot}$, in particular the constraints obtained from multi-messenger observations of GW170817 [67, 68, 69, 70, 71].
@@
The BNS-merger rate is described by the star-formation rate from Ref. [72] convolved with a distribution of delay times $t_{\mathrm{d}}$ that follows a power-law distribution $P(t_{\mathrm{d}})\propto t_{\mathrm{d}}^{-1}$ [73]. Using this model, we construct a catalogue containing about $N=3.8\times~{}10^{5}$ BNS that merge in the entire Universe during one year, consistent with other studies [74] (see Appendix B and Ref. [55]).
As we focus on the background, we must consider the flux contributions from non-detectable sources only. The detectability of a GW signal is estimated by calculating its SNR that, for a single GW detector, is given by
$\mathrm{SNR}^{2}:=4\int^{\infty}_{0}\frac{|\tilde{h}(f)|^{2}}{P_{\mathrm{n}}(f)}\,\mathrm{d}f\,,$ (7)
where $P_{\mathrm{n}}(f)$ is the noise PSD of the detector. For a network of $N_{\mathrm{det}}$ detectors, and assuming an uncorrelated noise among them, the SNR of a GW signal is calculated by summing the individual-detector SNRs in quadrature; we have here employed the GWFish package [75] to compute the SNR.
@@
We assume that individual resolvable sources can be identified when their SNR exceeds a threshold $\mathrm{SNR}_{\mathrm{thr}}$, so that the contribution from these resolvable sources is removed by subtracting each individual signal, however the process of parameter estimation of these sources inevitably comes with an uncertainty, so that the subtraction process is inherently affected and may introduce residual contributions to the background, as discussed in Ref. [76]. In this study, which represents a first investigation of the postmerger contribution, we assume that the subtraction is noiseless, and leave more realistic approaches for future work [77, 78, 79].
Following Refs. [80, 74], we take as SNR threshold for the detection of BNS mergers $\mathrm{SNR}_{\mathrm{thr}}=12$, but remark that for 3G detectors this value will depend on the details of the analysis pipeline and may affect the efficiency of signal subtraction. Indeed, it will be difficult to accurately measure the parameters of low-SNR sources, whose subtraction will introduce more noise in the SGWB compared to high-SNR sources. In this work, we investigate the following values for $\mathrm{SNR}_{\mathrm{thr}}=12,20,40$, and $80$.
III Results
@@
Figure 3 shows the SGWB from BNS mergers computed with our two descriptions for the modelling of the signal and detailed in Sec. II.1. More specifically, in the inspiral+prompt collapse model, hereafter “IPC” (blue-shaded areas), which is the one adopted so far in describing the SGWB from BNS mergers, the inspiral is assumed to be followed by a prompt collapse to a BH of the merger remnant. On the other hand, in the novel inspiral+postmerger model, hereafter “IPM”, proposed here (red-shaded areas), the signal includes the postmerger emission from a metastable HMNS that eventually collapses into a BH. For both models, we use ${\rm SNR}_{\mathrm{thr}}=12$, and the background was calculated using $10^{3}$ randomly sampled EOSs using the methodology described in Sec. II.2. The shaded regions correspond to the $68$ and $95$ percentiles of the resulting distribution, respectively, while the solid lines indicate the median values in the distributions.
@@
Figure 3: Energy density of the SGWB from BNS mergers, corresponding to either the IPC (blue-shaded areas) or the IPM (red-shaded areas) scenarios considered here. Shown with different transparencies are the $68$ and $95$ percentile contours, while the solid lines report the medians. Also reported are the power-integrated sensitivity curves of a single ET (dashed lines) and of a network composed of an ET detector and of two CE detectors (solid lines) relative to $1$ (grey lines) and $10$ (black lines) years of integration, respectively.
Figure 4: Left panel: Energy density of the SGWB from BNS mergers in the IPM scenario. Shown with different transparencies are the $68,95$, and $99$ percentile contours, while the solid lines report the medians. Different colours are used to show how $\Omega_{\rm GW}$ changes when assuming threshold SNR values of ${\rm SNR_{thr}}=12,20,40$, and $80$; note that $\tau_{2}=50\,{\rm ms}$ in all cases. Also reported are the power-integrated sensitivity curves of a single ET (dashed lines) and of a network composed of an ET detector and of two CE detectors (solid lines) relative to $1$ (grey lines) and $10$ (black lines) years of integration, respectively. Right panel: The same as in the left but when considering different values for the decay timescale $\tau_{2}=20,50$ and $250\,{\rm ms}$ (blue, green, and red, respectively). Note that ${\rm SNR_{thr}}=12$ in all cases.

### v2

@@ Abstract
The stochastic gravitational-wave background (SGWB) generated by the inspiral and merger of binary neutron stars is traditionally modelled assuming that the inspiral is promptly followed by the collapse of the merger remnant to a rotating black hole. While this is reasonable for the most massive binaries, it is not what is expected in general, as the remnant may survive for up to hundreds of milliseconds and radiate an amount of energy that is significantly larger than that lost during the whole inspiral. To account for this additional contribution to the SGWB, we consider a waveform model that includes both the inspiral and the postmerger emission. We show for the first time that for a large set of parameterized equations of state (EOSs) compatible with observational constraints, there is considerable spectral power in the $1-2\,{\rm kHz}$ range, distinct from that associated with the inspiral and leading to a dimensionless GW energy density $\Omega_{\rm GW}\simeq 10^{-10}-10^{-9}$. We discuss the enhanced detectability of the SGWB by third-generation detectors such as the Einstein Telescope and Cosmic Explorer, and show how it depends on the signal-to-noise ratio of foreground binaries and on the remnant lifetime. Interestingly, even a non-detection of the high-frequency part of the SGWB could provide valuable constraints on the remnant lifetime, offering novel insights into the postmerger dynamics and the EOS of nuclear matter.
@@
Introduction— The discovery of the binary neutron star (BNS) merger GW170817 LIGO and Virgo collaboration (2017a) marked a major milestone in multi-messenger astrophysics, as the observations of gravitational waves (GW) by LIGO-Virgo LIGO collaboration (2015); Acernese, F. et al. (2015) was followed by the detection of electromagnetic counterparts across multiple wavelengths Coulter, D. A. et al. (2017); Goldstein, A. et al. (2017); Savchenko, V. et al. (2017); Mooley et al. (2018); Ghirlanda, G. et al. (2019). These observations confirmed the connection between short gamma-ray bursts, kilonovae and BNS mergers and provided important information on the geometry of the jet launched after the merger as well as on the ejected neutron-rich material. Joint analysis of electromagnetic and GW data led to tight constraints of General Relativity collaboration (2019), whereas the identification of the host galaxy allowed an independent measurement of the Hubble parameter (see, e.g., LIGO and Virgo collaboration (2017b)). Combined with the second, GW-only, BNS detection collaboration (2020), the local merger rate is constrained to $10-1700\,{\rm Gpc}^{-3}\,{\rm yr}^{-2}$ Abbott, R. et al. (2023). This information is crucial for building population models and understanding the formation of binary compact objects.
@@
Observations of BNS mergers also provide a unique probe of the physics of extremely dense matter and allow to study the equations of state (EOS) of neutron star interiors. Since the EOS strongly affects the gravitational waveform, it leaves an imprint in the tidal deformation that takes place in the late inspiral stages and allows measuring the tidal deformability Bernuzzi et al. (2015); Hinderer, T. et al. (2016); Radice et al. (2014); Dietrich, Tim et al. (2019); Dietrich et al. (2019). In the case of the GW170817 event, GW observations allowed to constrain the tidal deformability parameter to $\tilde{\Lambda}\lesssim 700$, where $\tilde{\Lambda}$ is the mass-weighted combination of the dimensionless deformability parameters Abbott, B. P. et al. (2018a); LIGO and Virgo collaboration (2019).
@@
Additional and complementary information on the EOS can be extracted from the postmerger signal emitted by the hypermassive neutron star (HMNS) formed after the neutron stars have coalesced Shibata et al. (2017); Baiotti and Rezzolla (2017); Radice et al. (2020). In particular, a number of recent studies have shown that the GW emission from the HMNS is concentrated at specific frequencies, resulting in relatively narrow spectral features Stergioulas et al. (2011); Bauswein and Janka (2012); Takami et al. (2014a); Bauswein and Stergioulas (2015a); Rezzolla and Takami (2016); Takami et al. (2015a); Topolski et al. (2024). The postmerger signal has not been detected so far, as the relevant frequencies are above those observable with current ground-based detectors. However, future third-generation (3G) detectors, such as the Einstein Telescope (ET) Maggiore, M. et al. (2020) and the Cosmic Explorer (CE) Evans, M. et al. (2021) will be able to detect and characterize the postmerger signal for sufficiently nearby systems.
@@
An alternative method to study populations of compact binaries is by observing the stochastic GW background (SGWB) created by the incoherent superposition of unresolved sources Regimbau (2011); Christensen (2019). Different types of sources can contribute to the SGWB, including compact binaries, core-collapse supernovae and isolated spinning neutron stars (e.g. Buonanno et al., 2005; Regimbau and Mandic, 2008; Crocker et al., 2015). Observing the SGWB will allow probing high-redshift populations, which are otherwise not accessible to ground-based detectors, although various methods are being developed to extract the SGWB from the detector noise in data from Advanced LIGO, Advanced Virgo and Kagra (LVK) Regimbau (2022); Renzini et al. (2022). Current upper limits for ground-based detectors are $\Omega_{\rm{GW}}(f)\leq 3.4\times 10^{-9}$ at $f=25$ Hz, where $\Omega_{\rm{GW}}$ is the dimensionless energy density of the SGWB Abbott, B. P. et al. (2017, 2018b); Abbott, B. P. et al. (2019, 2021); Abbott, R. et al. (2021). These upper limits contain valuable information on high-redshift populations of compact binaries, complementary to that of individual detections (e.g. Kowalska-Leszczynska et al., 2015; Dvorkin et al., 2016; Nakazato et al., 2016; Périgois et al., 2021; Cusin et al., 2019; Callister et al., 2020; Périgois et al., 2022; Lehoucq et al., 2023). For the LVK detectors, the dominant contribution to the SGWB is expected to be that of binary black holes. However, this will not be the case for ET and CE, since almost all the binary black hole mergers will be detected individually and subtracted from the data. The remaining SGWB will thus contain information on the unresolved BNS mergers.
@@
In the case of the postmerger emission from a population of BNS mergers, we expect that if HMNS are formed, the characteristic GW frequencies will be enhanced in the SGWB due to the large number of unresolved sources contributing to the signal. This opens up the possibility to detect this EOS-dependent feature and to obtain complementary information on the postmerger stage. The goal of this Letter is to model the SGWB from a population of BNS, including the postmerger stage, and to highlight the existence of a so-far neglected additional and dominant component that can be detected with ET and CE.
@@
BNS inspiral and postmerger waveforms— To construct a hybrid waveform, we combine in the frequency domain the analytic prescriptions referring to the inspiral and the postmerger. This combination takes place at a frequency $f_{\rm tr}$, where we move from the inspiral to the postmerger description (see End Matter for details). This frequency must be significantly smaller than the main frequencies appearing in the power spectral density (PSD) of the postmerger signal, i.e., $f_{1}$ and $f_{2}$, where $f_{1}<f_{2}$, so that no power from the postmerger phase is missed. Hereafter, for simplicity we set $f_{\rm tr}=f_{1}/2$, but the overall results depend only weakly on this assumption. The PSD in the $f_{1}$ frequency is significantly suppressed in the case on unequal-mass binaries but is here included for consistency with the use of a population of neutron-star binaries with equal-mass components. In practice, because the corresponding power is about one order of magnitude smaller than that in the $f_{2}$ peak, its impact in shaping the high-frequency component of the SGWB is very small.
@@
The equation of state— In view of the uncertainties in the EOS and their effect on the waveform, we adopt an agnostic approach in this work. We use the findings of Refs. Altiparmak et al. (2022); Ecker and Rezzolla (2022, 2023); Jiang et al. (2023), who constructed more than $10^{7}$ EOSs consistent with constraints from fundamental physics and astronomical observations of neutron stars. Leveraging on this extensive dataset, it was possible to derive informative constraints on the relationship between the chirp mass $\mathcal{M}_{\mathrm{chirp}}$ and tidal deformability $\Lambda$ in terms of a probability density function (PDF) for these two quantities. A-priori, a single EOS does not correspond to a unique line in the $\mathcal{M}_{\mathrm{chirp}}$-$\Lambda$ plane because the same chirp mass may result from many combinations of BNS masses. However, when restricting the analysis to symmetric binaries, as we do throughout the rest of this work, this ambiguity disappears. In this case, for a given EOS, each chirp mass is uniquely associated to a single value of $\Lambda$. Therefore, extracting an EOS from a given distribution is equivalent to drawing a single value of $\Lambda$ from the PDF at each chirp mass. Similarly, extracting a certain value of $\Lambda$ from the allowed range is equivalent to singling-out a specific EOS.
Although the PDF found in Ref. Altiparmak et al. (2022) for a given chirp mass is not exactly a Gaussian, it is so to a first approximation [see Fig. (4) in Altiparmak et al. (2022)]. Hence, when assuming an underlining Gaussian distribution, we can relate the tidal deformability and the chirp mass (in solar masses) via a simple power-law of the type $\Lambda_{p}=\alpha+\beta\mathcal{M}^{\gamma}_{\mathrm{chirp}}$ where $\Lambda_{p}$ is normally distributed between the minimum and maximum values given by Altiparmak et al. (2022) $\Lambda_{\rm min\,(max)}=a+b\,\mathcal{M}_{\rm chirp}^{c}$ with $a=-50\,(-20)$, $b=500\,(1800)$, and $c=-4.5\,(-5.0)$. We here consider $\gamma=-3$, but we have checked that varying $\gamma$ leads to very little differences in SGWB, as the mass distribution of BNS has a clear main peak. In this way, we can build a set of fictitious $10^{3}$ EOSs such that the corresponding tidal deformability is normally distributed.
@@
Once an EOS is set, we can compute the inspiral waveform. For the postmerger waveform, we need to know the characteristic frequencies $f_{1}$ and $f_{2}$ given the selected EOS and hence tidal deformability. Luckily, this is rather straightforward to do since universal relations have been found that provide analytic fits (see, e.g., Bauswein and Janka (2012); Takami et al. (2014a); Bauswein and Stergioulas (2015a); Rezzolla and Takami (2016)). Here, we employ the expressions for $f_{1}$ and $f_{2}$ proposed in Refs. Rezzolla and Takami (2016); Topolski et al. (2024) (and also employed in Bose et al. (2018)). We note that being only “quasi”-universal relations, these expressions come with some intrinsic uncertainties as estimated from the numerical simulations and the corresponding fits (see discussion in Topolski et al. (2024)). These uncertainties, however, do not play an important role here as we are interested in a SGWB that is naturally broad given the much larger uncertainties in the EOS.
@@
Stochastic gravitational-wave background— The energy density in GWs per logarithmic unit of frequency and in units of the critical density of the Universe, $\rho_{\mathrm{c}}=3H_{0}/8\pi G$, is given by $\Omega_{\mathrm{GW}}:=({1}/{\rho_{c}})d\rho_{\mathrm{GW}}/{d}\ln{f}=({f}/{c\,\rho_{\mathrm{c}}})F(f)$, where $\rho_{\mathrm{GW}}$ is the GW energy density and $F(f)$ is the energy flux. When considering a population of $N$ sources, the energy flux per unit frequency is given by Allen and Romano (1999): $F(f):=({\pi c^{3}\,f^{2}}/{2G\,T})\sum_{i=1}^{N}[|\tilde{h}_{i}^{+}(f)|^{2}+|\tilde{h}_{i}^{\times}(f)|^{2}]$, where the index $i$ sums over the $N$ sources and $\tilde{h}_{i}^{+}(f)$ and $\tilde{h}_{i}^{\times}(f)$ are the Fourier transforms of the two GW polarization modes. The number of sources is taken to be proportional to the total observation time $T$. We have checked that our calculation converges at $T=1$ year and will use this value in the following (note that the background is stationary).
@@
In order to construct our BNS catalogue, we follow the baseline_delays model from Ref. Lehoucq et al. (2023) and, more specifically, we assume all binaries are symmetric in mass and follow a Gaussian mass distribution with a mean of $1.33\,M_{\odot}$ and a standard deviation of $0.09\,M_{\odot}$. The BNS-merger rate is described by the star-formation rate from Ref. Vangioni et al. (2015) convolved with a distribution of delay times $t_{\mathrm{d}}$ that follows a power-law distribution $P(t_{\mathrm{d}})\propto t_{\mathrm{d}}^{-1}$ Chruslinska et al. (2018). Using this model, we construct a catalogue containing about $N=3.8\times~{}10^{5}$ BNS that merge in the entire Universe during one year, consistent with other studies Iacovelli et al. (2022) (see Ref. Lehoucq et al. (2023)).
@@
As we focus on the background, we must consider the flux contributions from non-detectable sources only. We assume that individual resolvable sources can be identified when their signal-to-noise ratio (SNR) exceeds a threshold $\mathrm{SNR}_{\mathrm{thr}}=12$  Meacher et al. (2015); Iacovelli et al. (2022), so that the contribution from these resolvable sources is removed by subtracting each individual signal, and we use the GWFish package Dupletsa et al. (2023) to compute the SNR. In this initial study we assume that the subtraction is noiseless, and leave more realistic approaches for future work Sachdev et al. (2020a); Regimbau et al. (2017); Sachdev et al. (2020b); Zhou et al. (2023).
@@
Results— Figure 1 shows the SGWB from BNS mergers computed with our two descriptions for the modelling of the signal. More specifically, in the inspiral+prompt collapse model, hereafter “IPC” (blue-shaded areas), which is the one adopted so far in describing the SGWB from BNS mergers, the inspiral is assumed to be followed by a prompt collapse to a BH of the merger remnant. On the other hand, in the novel inspiral+postmerger model, hereafter “IPM”, proposed here (red-shaded areas), the signal includes the postmerger emission from a metastable HMNS that eventually collapses into a BH. For both models, we use ${\rm SNR}_{\mathrm{thr}}=12$, and the background was calculated using $10^{3}$ randomly sampled EOSs using the methodology described above. The shaded regions correspond to the $68$ and $95$ percentiles of the resulting distribution, respectively, while the solid lines indicate the median values in the distributions.
@@
Figure 1: Energy density of the SGWB from BNS mergers, corresponding to either the IPC (blue-shaded areas) or the IPM (red-shaded areas) scenarios. Shown with different transparencies are the $68$ and $95$ percentile contours, while the solid lines report the medians. Also outlined are the power-integrated sensitivity curves of a single ET (dashed lines) and of a network composed of an ET detector and of two CE detectors (solid lines) relative to $1$ (grey lines) and $10$ (black lines) years of integration, respectively.
Figure 2: Top panel: The same as in Fig. 1 but where different colours show how $\Omega_{\rm GW}$ changes when assuming threshold SNR values of ${\rm SNR_{thr}}=12,20,40$, and $80$; note that $\tau_{2}=50\,{\rm ms}$ in all cases. Bottom panel: The same as in the top but when considering different values for the decay timescale $\tau_{2}=20,50$ and $250\,{\rm ms}$ (blue, green, and red, respectively). Note that ${\rm SNR_{thr}}=12$ in all cases.